# A Systematic Study of Cross-Layer KV Sharing for Efficient LLM Inference

### v1

@@ Abstract
Recently, sharing key-value (KV) cache across layers has been found effective in efficient inference of large language models (LLMs). To systematically investigate different techniques of cross-layer KV sharing, we propose a unified framework that covers several recent methods and their novel variants. We conduct comprehensive experiments on all the configurations of the framework, evaluating their generation throughput and performance in language modeling and downstream tasks. We find that when reducing the size of the KV cache by 2×2$\times$2 ×, most configurations can achieve competitive performance to and higher throughput than standard transformers, but when further reducing the size of the KV cache, pairing queries of all layers with KVs of upper layers can better maintain performance, although it also introduces additional training cost and prefilling latency. We hope that this work will help users choose the appropriate approach according to their requirements and facilitate research on the acceleration of LLM inference.
@@ 1 Introduction
In this study, we consider a unified framework for cross-layer KV sharing, of which LCKV, CLA, and YOCO can be seen as special configurations. We then empirically test all the configurations of the framework, including several novel ones that have never been considered in previous work. Our experiments show that all the configurations can achieve significantly higher throughput than the standard transformer when the prompt is short, but the throughput of the configurations that compute the KVs at the top layers degrades dramatically when the prompt is long. On the other hand, while the performance of most configurations is comparable with that of the standard transformer when only half of the layers rely on the KVs computed by the other layers, it is the configurations that compute the KVs at bottom layers whose performance degrades the most when more layers become reliant on the other layers for the KVs. We hope our framework and empirical studies would help users interested in cross-layer KV sharing to choose methods and configurations according to their throughput and performance requirements. Our code is available at https://github.com/whyNLP/LCKV.
@@ 2 Existing Methods
You Only Cache Once (YOCO) Sun et al. (2024) computes the KVs of only the top layer of the transformer, which are paired with the queries of the top-half of the layers. The bottom-half of the layers uses efficient attention to achieve a constant cache size. Goldstein et al. (2024) uses a similar sharing pattern to YOCO, but further compresses the size of the KV cache.
Cross-Layer Attention (CLA) Brandon et al. (2024) uniformly divides transformer layers into multiple groups of adjacent layers. In each group, it pairs the queries of all the layers with the KVs of the bottom layer. Zuhri et al. (2024) shares the KVs in the same way as CLA, but applies a more efficient training scheme. Liu et al. (2024) groups every two adjacent layers in the middle-to-deep portion and compresses the KV cache across each group. Chen et al. (2024) groups non-adjacent layers and pairs the queries of the upper layer with the KVs of the lower layer in each group. Rajput et al. (2024) uses a combination of the sliding window attention layers and a similar sharing pattern to CLA. Liao and Vargas (2024); Mu et al. (2024); Rajabzadeh et al. (2024) apply the sharing pattern similar to CLA to the computed attention weights instead of KVs.
@@ 3 A Unified Framework
We define a configuration by partitioning transformer layers and positioning target layer(s) differently. We choose the layer partitioning from { pizza, sandwich, lasagna } and choose the target layer positioning from { bottom, top, middle } [^1] We also consider positioning at quarter and three-quarter, which is discussed in Appendix D.. The pizza partitioning sets the first $l-1$ layers as KV layers. The sandwich partitioning sets the first $\lceil\frac{l-1}{2}\rceil$ layers and the last $\lfloor\frac{l-1}{2}\rfloor$ layers as KV layers. For the remaining $L-l+1$ consecutive layers in both pizza and sandwich, their target layer is positioned at either the top, the middle, or the bottom of these layers. The lasagna partitioning uniformly divides the $L$ layers into $l$ groups of consecutive layers. For each group except the first, the target layer of all the layers within the group is positioned at either the top, the middle, or the bottom of these layers. For the first group, however, we always set the bottom layer as the target layer because we empirically find that there is a significant drop in performance if the first layer is not a KV layer.
@@
Table 1: All the configurations in our unified framework for cross-layer KV sharing. Red layers are KV layers. Each arrow points to a target layer from the layers whose queries are paired with its KV.
@@ 4.1 Generation Throughput
Figure 1(a) reports the maximum throughput. When the prompt is short (i.e., 5+2043), the prefilling time can be ignored and the generation throughputs of all the nine configurations are almost identical, which are much higher than the baseline throughput and increase as the number of KV layers decreases. When the prompt is long (i.e., 512+1024), the prefilling time becomes significant for the top and middle configurations because of iterative encoding of the prompt. Consequently, their throughputs degrade dramatically, falling below the baseline in some cases. On the other hand, the bottom configurations still achieve significantly higher throughputs than the baseline because no additional computation for prompt is required.
@@ 4.2 Performance on Small Training Set
We train the standard Llama and the nine configurations with 110M and 1.1B parameters from scratch [^3] We also tried model initialization with pre-trained models, the results of which are shown in Appendix C. on the Minipile dataset Kaddour (2023) with 1.7B tokens for one epoch and two epochs, respectively, and evaluate their perplexity. The training details are shown in Appendix A.
@@ 4.3 Performance on Large Training Set
Figure 1(c) reports the perplexity and average accuracy of downstream tasks. Detailed results of downstream tasks are shown in Appendix B. It can be seen that the sandwich-top configuration
@@
performs better than the two bottom configurations in both perplexity and downstream task accuracy, except for an outlier of the lasagna-bottom configuration with 7 KV layers in downstream task accuracy. The sandwich-middle configuration performs best when the number of KV layers is small.
@@ Appendix A Model and Training Details
Appendix B Detailed Downstream Task Results
@@
Appendix C Initializing with Pre-trained Models
@@
Appendix D More Options for Target Layer Positioning

### v2

@@ Abstract
Recently, sharing key-value (KV) cache across layers has been found effective in efficient inference of large language models (LLMs). To systematically investigate different techniques of cross-layer KV sharing, we propose a unified framework that covers several recent methods and their novel variants. We conduct comprehensive experiments on all the configurations of the framework, evaluating their generation throughput and performance in language modeling and downstream tasks. We find that when reducing the size of the KV cache by 2×2$\times$2 ×, most configurations can achieve higher throughput than standard transformers while maintaining competitive performance. When further reducing the size of the KV cache, however, pairing queries of all layers with KVs of upper layers performs better, at the expense of additional training cost and prefilling latency. We hope that this work will help users make more informed choices of cross-layer KV sharing approaches and facilitate future research on efficient LLM inference.
@@ 1 Introduction
In this study, we consider a unified framework for cross-layer KV sharing, of which LCKV, CLA, and YOCO can be seen as special configurations. We then empirically test all the configurations of the framework, including several novel ones that have never been considered in previous work. Our experiments show that, with respect to throughput, all the configurations can achieve significantly higher throughput than the standard transformer when the prompt is short; but when the prompt is long, the throughput of the configurations that compute the KVs at the top layers degrades dramatically. With respect to performance, when only half of the layers rely on the KVs computed by the other layers, the performance of most configurations is comparable with that of the standard transformer; when more layers become reliant on the other layers for the KVs, the configurations that compute the KVs at the bottom layers suffer the greatest performance degradation. We hope our framework and empirical studies would help users interested in cross-layer KV sharing to make more informed choices of methods and configurations according to their throughput and performance requirements. Our code is available at https://github.com/whyNLP/LCKV.
@@ 2 Existing Methods
You Only Cache Once (YOCO) Sun et al. (2024) computes the KVs of only the middle layer of the transformer, which are paired with the queries of the top-half of the layers. The bottom-half of the layers uses efficient attention to achieve a constant cache size. Goldstein et al. (2024) uses a similar sharing pattern to YOCO, but further compresses the size of the KV cache.
Cross-Layer Attention (CLA) Brandon et al. (2024) uniformly divides transformer layers into multiple groups of adjacent layers. In each group, it pairs the queries of all the layers with the KVs of the bottom layer. Zuhri et al. (2024) shares the KVs in the same way as CLA, but applies a more efficient training scheme. Liu et al. (2024) groups every two adjacent layers in the middle-to-deep portion and compresses the KV cache in each group. Chen et al. (2024) groups non-adjacent layers and pairs the queries of the upper layer with the KVs of the lower layer in each group. Rajput et al. (2024) uses a combination of the sliding window attention and a sharing pattern similar to CLA. Liao and Vargas (2024); Mu et al. (2024); Rajabzadeh et al. (2024) apply sharing patterns similar to CLA to the computed attention weights instead of KVs.
@@ 3 A Unified Framework
We define a configuration by partitioning transformer layers and positioning target layer(s) differently. We choose the layer partitioning from { pizza, sandwich, lasagna } and choose the target layer positioning from { bottom, top, middle } [^1] We also consider positioning at quarter and three-quarter, which is discussed in Appendix E.. The pizza partitioning sets the first $l-1$ layers as KV layers. The sandwich partitioning sets the first $\lceil\frac{l-1}{2}\rceil$ layers and the last $\lfloor\frac{l-1}{2}\rfloor$ layers as KV layers. For the remaining $L-l+1$ consecutive layers in both pizza and sandwich, their target layer is positioned at either the top, the middle, or the bottom of these layers. The lasagna partitioning uniformly divides the $L$ layers into $l$ groups of consecutive layers. For each group except the first, the target layer of all the layers within the group is positioned at either the top, the middle, or the bottom of these layers. For the first group, however, we always set the bottom layer as the target layer because we empirically find that there is a significant drop in performance if the first layer is not a KV layer.
@@
Table 1: All the configurations in our unified framework for cross-layer KV sharing. Red layers are KV layers. Each arrow points to a target layer from the layers whose queries are paired with its KV. The sandwich-top configuration corresponds to LCKV, the pizza-bottom configuration corresponds to YOCO, and the lasagna-bottom configuration corresponds to CLA.
@@ 4.1 Generation Throughput
Figure 1(a) reports the maximum throughput [^3] The throughput at different batch sizes is shown in Appendix B.. When the prompt is short (i.e., 5+2043), the prefilling time can be ignored and the generation throughputs of all the nine configurations are almost identical, which are much higher than the baseline throughput and increase as the number of KV layers decreases. When the prompt is long (i.e., 512+1024), the prefilling time becomes significant for the top and middle configurations because of iterative encoding of the prompt. Consequently, their throughputs degrade dramatically, falling below the baseline in some cases. On the other hand, the bottom configurations still achieve significantly higher throughputs than the baseline because no additional computation for prompt is required.
@@ 4.2 Performance on Small Training Set
We train the standard Llama and the nine configurations with 110M and 1.1B parameters from scratch [^4] We also tried model initialization with pre-trained models, the results of which are shown in Appendix D. on the Minipile dataset Kaddour (2023) with 1.7B tokens for one epoch and two epochs, respectively, and evaluate their perplexity. The training details are shown in Appendix A.
@@ 4.3 Performance on Large Training Set
Figure 1(c) reports the perplexity and average accuracy of downstream tasks.
@@
Detailed results of downstream tasks are shown in Appendix C. It can be seen that the sandwich-top configuration performs better than the two bottom configurations in both perplexity and downstream task accuracy, except for an outlier of the lasagna-bottom configuration with 7 KV layers in downstream task accuracy. The sandwich-middle configuration performs best when the number of KV layers is small.
@@ 5 Conclusion
Acknowledgements
This work was supported by HPC Platform of ShanghaiTech University.
@@ Appendix A Model and Training Details
Appendix B Throughput at Different Batch Sizes
Figure 2 reports the generation throughput of the standard Llama and the nine configurations with different numbers of KV layers at different batch sizes. The highest point of each curve indicates the maximum throughput of the model, which has been shown in Figure 1(a), and the rightmost point indicates the maximum batch size. It can be seen that, at any given batch size, the throughput of the nine configurations is higher than the baseline throughput and increases as the number of KV layers decreases.
Figure 2: Throughput of 1.1B models at different batch sizes on an RTX 3090 (24GB) GPU with a prompt length of 5 and a generation length of 2043.
Appendix C Detailed Downstream Task Results
@@
Appendix D Initializing with Pre-trained Models
@@
Appendix E More Options for Target Layer Positioning